Solve boolean algebra rule: A+ĀB=A+B.I want its explaination how it can be derived... I mean L.H.S=R.H.S

Grade:12th pass

1 Answers

Vaidehi
6 years ago
by distributive law A+(B.C)=(A+B).(B+C) 
apply the same here Ā).(A+B)=1.(A+B)=(A+B) ;(Since A+Ā=1)A+ĀB=(A+